OpenStudy (sh3lsh):

Linear Alg. Give a 2x2 matrix A where this property holds true. 1) A^2 = A, all entries are nonzero 2) A ^3 = A, all entries are nonzero For both, I can use the orthogonal projections, but when could I use the reflection matrix?

OpenStudy (anonymous):

if \(A\) is a reflection, then \(A^2=AA=I\) so it would not satisfy 1. however, \(A^3= AA^2=A\) would be satisfied

OpenStudy (sh3lsh):

Again, you're the best.